OpenStudy (anonymous):

HELP ME PLEASE , ILL MEDAL At a Phillies game, they sell a small cup of soda for $2.50 and a large cup of soda for $5. The booth that sold the drinks collected a total of $7000 from the sale of soda. If they sold a total of 1600 cups of soda. How many small cups did they sell?

OpenStudy (texaschic101):

S = # of small cups L = # of large cups total of 1600 cups of soda were sold (both small and large) S + L = 1600 $7000 was collected ($ 2.50 for S and $ 5 for L) 2.50S + 5L = 7000 now that you have the 2 equations, we can now solve this by substitution. S + L = 1600 ----> L = 1600 - S now just sub in 1600 - S in for L back into the other equation and solve for S, the number of small sodas sold. Let me start you off..... L = 1600 - S 2.50S + 5L = 7000 2.50S + 5(1600 - S) = 7000 now just distribute through the parenthesis and solve for S.